The Jones Company has just completed the third year of a​ five-year MACRS recovery period for a piece of equipment it originally purchased for $ 299,000.

a. What is the book value of the​ equipment?

b. If Jones sells the equipment today for $ 175,000 and its tax rate is 35 %​, what is the​ after-tax cash flow from selling​ it?

c. Just before it is about to sell the​ equipment, Jones receives a new order. It can take the new order if it keeps the old equipment. Is there a cost to taking the order and if​ so, what is​ it? Explain.​ (Assume the new order will consume the remainder of the​ machine's useful​ life.) Note​: Assume that the equipment is put into use in year 1.
Answer #1

a.Book value of Equipment = Original cost - Depreciation charged for 3 years under MACRS

= 299000*(100-20-32-19.20)%

= $86,112

b.Selling price = $175000

Gain on Sale = 175000-86112 = $88,888

Tax on Gain = 88,888*35% = $31,110.8

After tax Sales price = 175000 - 31,110.8 = $143,889.2

c.The cost is the opportunity cost of selling machine i.e. after tax sales price of $143,889.2